OpenStudy (anonymous):

Sam picked a card from a standard deck. What is the probability that Sam picked a heart or a king? one over thirteen sixteen over fifty two seventeen over fifty two sixteen over fifty three

OpenStudy (amistre64):

so, how many cards are possible in the event? how many hearts/kings are there in total?

OpenStudy (anonymous):

i know theres 52 cards in a deck, and 4 cards of each number. and 4 suits. i just cant remember the number of cards in each suit.

OpenStudy (amistre64):

52/4 cards in each suit

OpenStudy (anonymous):

Ohh 13.

OpenStudy (amistre64):

so, 13 hearts, and 4 kings ... but we already count a king in the hearts so we cant reasonable count it again

OpenStudy (anonymous):

So, do i do 26/52?

OpenStudy (amistre64):

16, but yeah

OpenStudy (anonymous):

Oh. So its the third one.

OpenStudy (amistre64):

17/52 is not 16/52 you should really write your options in mathical notation

OpenStudy (amistre64):

a formula they use is:\[P(A)+P(B)-P(A\cap B)\]

OpenStudy (anonymous):

i meant the second. i dont know why i said third ;c

OpenStudy (amistre64):

\[\frac{13}{52}+\frac{4}{52}-\frac{1}{52}\]

OpenStudy (anonymous):

\[\frac{ 16 }{ 52 }\]

OpenStudy (amistre64):

good

OpenStudy (anonymous):

Thanks!